# Understanding the Performance Gap in Preference Learning: A Dichotomy of RLHF and DPO

### v2

@@ Abstract
We present a fine-grained theoretical analysis of the performance gap between reinforcement learning from human feedback (RLHF) and direct preference optimization (DPO) under a representation gap. Our study decomposes this gap into two sources: an explicit representation gap under exact optimization and an implicit representation gap under finite samples. In the exact optimization setting, we characterize how the relative capacities of the reward and policy model classes influence the final policy qualities. We show that RLHF, DPO, or online DPO can outperform one another depending on type of model mis-specifications. Notably, online DPO can outperform both RLHF and standard DPO when the reward and policy model classes are isomorphic and both mis-specified. In the approximate optimization setting, we provide a concrete construction where the ground-truth reward is implicitly sparse and show that RLHF requires significantly fewer samples than DPO to recover an effective reward model—highlighting a statistical advantage of two-stage learning. Together, these results provide a comprehensive understanding of the performance gap between RLHF and DPO under various settings, and offer practical insights into when each method is preferred.
@@ 1 Introduction
However, the assumptions of tabular parameterization and realizability often do not hold in practice, particularly when the reward model is significantly smaller than the policy model (e.g., 66B vs. 175175B in Ouyang et al. (2022), indicating a clear disparity in representational capacity), when the policy model class is heavily restricted due to limited computational resources, or when the reward model is sub-optimal owing to limited preference data. These situations are examples of model mis-specification, a common issue in practice due to limitations in model capacity or data. Consequently, one should not expect DPO to perform identically to RLHF under model mis-specifications. This motivates the central question of our investigation:
@@
To quantify the problem, we choose the performance metric as the expected value of the original regularized bandit problem using the ground-truth reward $r^{\star}$ ($x$ is a prompt, and $y$ is a response): $V_{r^{\star}}^{\pi}:=\mathbb{E}_{x\sim\rho}\left[\mathbb{E}_{y\sim\pi(\cdot|x)}[r^{\star}(x,y)]-\beta\mathsf{KL}\left(\pi(\cdot|x)\|\pi_{{\mathsf{ref}}}(\cdot|x)\right)\right]$, where $\rho$ is a pre-fixed distribution over prompts, $\pi$ is a distribution over responses, and $\pi_{{\mathsf{ref}}}$ is the base model. Let $\pi^{\star}:=\operatorname{argmax}_{\pi}V_{r^{\star}}^{\pi}$ be the ideal optimal policy.
Our contributions. We study the performance differences between two-stage RLHF and DPO under a representation gap, from an optimization perspective. Our contributions are listed as follows:
Sec. 3.1 $V_{r^{\star}}^{\pi_{\textup{RLHF}}}=V_{r^{\star}}^{\pi_{\textup{DPO}}}=V_{r^{\star}}^{\pi^{\star}}$ $\mathcal{L}_{\textup{DPO}}^{\textup{online}}(\pi(\cdot|x))\propto-V_{r^{\star}}^{\pi(\cdot|x)}+\mathcal{O}(\delta^{2})$ Sec. 3.3 $V_{r^{\star}}^{\pi_{\textup{RLHF}}}\leq V_{r^{\star}}^{\pi_{\textup{DPO}}}$ Sec. 3.2 $V_{r^{\star}}^{\pi_{\textup{RLHF}}}\geq V_{r^{\star}}^{\pi_{\textup{DPO}}}$ $V_{r^{\star}}^{\pi_{\textup{RLHF}}}\geq V_{r^{\star}}^{\pi_{\textup{DPO}}^{\textup{online}}}$ Sec. 3.4 It depends on
@@
[…]
@@
$\bullet$ When assuming exact optimization, i.e., optimization with infinite data, we study the fine-grained representation gap under different settings of reward and policy class mis-specifications in Section 3. Main results are visualized in Figure 1.
No model mis-specification: We show that the RLHF and DPO policies both achieve the performance of $\pi^{\star}$, and online DPO can further close the gap between optimization paths.
Policy model mis-specification: We show that the RLHF policy is still optimal under the model class, while the DPO policy can be sub-optimal, and online DPO cannot bridge the gap.
Reward model mis-specification: We show that the DPO policy is still optimal, while the RLHF policy can be sub-optimal due to learning based on a sub-optimal reward model.
Double model mis-specification: When policy and reward model classes are isomorphic, then they should have identical performance, while online DPO can outperform both of them. Otherwise, there is no consistent performance gap, and the comparison result depends on the qualities of (surrogate) reward models. We also give a preliminary guide for reward learning under mis-specifications.
$\bullet$ For approximate optimization, i.e., the finite-sample regime, we study the implicit representation gap incurred by statistical efficiencies in Section 4. We construct a simple task where the ground-truth reward to is a dual-token linear function with feature dimension $d$ and implicit sparsity $k$, and the total number of samples is $n$. Even without mis-specifications, we can reveal a separation between RLHF and DPO under this setting: the estimation error of DPO is $\Omega(d/n)$, while reward learning in RLHF can effectively leverage sparsity, decreasing the error to $\tilde{\mathcal{O}}(\sqrt{k\log d/n})$. This result indicates that DPO is less data-efficient than RLHF, leading to inferior performance.
Finally, we conduct numerical experiments to corroborate these theoretical findings in Section 5.
@@ 2 Preliminaries
Notation. Let $\sigma:\mathbb{R}\rightarrow\mathbb{R}$ be the sigmoid function, where $\sigma(x)=1/(1+\exp(-x))$. For any set $\mathcal{X}$, $\Delta(\mathcal{X})$ represents the set of probability distributions over $\mathcal{X}$. $\mathsf{sg}\left(\right)$ is the stopping-gradient operator, where $\nabla_{\theta}[\mathsf{sg}\left(f(\theta)\right)]=\boldsymbol{0}$. Let $e_{k}$ be a one-hot vector with $1$ on its $k\textsuperscript{th}$ entry and 0 on other entries. We use $f(\theta){\stackrel{{\scriptstyle\scriptstyle\nabla}}{{=}}}g(\theta)$ to indicate $\nabla_{\theta}f(\theta)=\nabla_{\theta}g(\theta)$.
Bandits and Policies. A bandit is defined by a state space $\mathcal{X}$, an action space $\mathcal{Y}$, and a reward function $r:\mathcal{X}\times\mathcal{Y}\to\mathbb{R}$. A policy $\pi:\mathcal{X}\rightarrow\Delta(\mathcal{Y})$ represents a probability distribution over actions given a state. Note that, we sometimes omit the prompt $x$ for simplicity, so that $\pi\in\Delta(\mathcal{Y})$.
@@
$V_{r}^{\pi}:=\underset{x\sim\rho}{\mathbb{E}}\left[\underset{y\sim\pi(\cdot|x)}{\mathbb{E}}[r(x,y)]-\beta\,\mathsf{KL}\left(\pi(\cdot|x)\|\pi_{{\mathsf{ref}}}(\cdot|x)\right)\right]\;,$
where $\beta>0$ is the regularization coefficient, and $\pi_{{\mathsf{ref}}}$ is a fixed reference policy. Let $r^{\star}$ denote the ground-truth reward function, and $\pi^{\star}$ denote the optimal policy for $V_{r^{\star}}^{\pi}$. A well-known fact (Rafailov et al., 2023) is that $\pi^{\star}(y|x)=\pi_{{\mathsf{ref}}}(y|x)\exp(r^{\star}(x,y)/\beta)/Z(x)$, where $Z(x):=\sum_{y\in\mathcal{Y}}\pi_{{\mathsf{ref}}}(y|x)\exp(r^{\star}(x,y)/\beta)$ is the partition function. The goal of preference-based policy learning is to find a policy $\pi_{\theta}\in\Pi$ that maximizes $V_{r^{\star}}^{\pi_{\theta}}$. We define the oracle value as $V^{\star}_{\Pi}:=\max_{\pi\in\Pi}V_{r^{\star}}^{\pi}$.
@@
Two-stage approach of RLHF. RLHF proceeds in two stages. First, the reward learning stage finds a reward model $r_{\textup{RLHF}}\in\mathcal{F}$ by maximizing the population MLE objective:
$r_{\textup{RLHF}}=\underset{r_{\phi}\in\mathcal{F}}{\operatorname{argmax}}\;\underset{x\sim\rho;y,y^{\prime}\sim\pi_{{\mathsf{ref}}}(\cdot|x)}{\mathbb{E}}\left[p^{\star}(y\succ y^{\prime}|x)\log\sigma(r_{\phi}(x,y)-r_{\phi}(x,y^{\prime}))+p^{\star}(y^{\prime}\succ y|x)\log\sigma(r_{\phi}(x,y^{\prime})-r_{\phi}(x,y))\right]\;.$
Direct approach of DPO. By leveraging the surrogate reward $\hat{r}_{\theta}(x,y):=\beta\log\frac{\pi_{\theta}(x,y)}{\pi_{{\mathsf{ref}}}(x,y)}$, DPO bypasses reward learning and directly learns the policy from preference data:
$\pi_{\text{DPO}}=\underset{\pi_{\theta}\in\Pi}{\operatorname{argmax}}\;\underset{x\sim\rho;y,y^{\prime}\sim\pi_{{\mathsf{ref}}}(\cdot|x)}{\mathbb{E}}\left[p^{\star}(y\succ y^{\prime}|x)\log\sigma\left(\hat{r}_{\theta}(x,y)-\hat{r}_{\theta}(x,y^{\prime})\right)+p^{\star}(y^{\prime}\succ y|x)\log\sigma\left(\hat{r}_{\theta}(x,y^{\prime})-\hat{r}_{\theta}(x,y)\right)\right]\;.$
We also consider an online variant of DPO, where the pairwise data are sampled from a distribution $\pi^{\mathsf{s}}$ which could depend on the current policy. Online DPO then minimizes the modified loss:
$\mathcal{L}_{\textup{DPO}}^{\textup{online}}(\pi_{\theta}(\cdot|x))=-\underset{y,y^{\prime}\sim\mathsf{sg}\left(\pi^{\mathsf{s}}(\cdot|x)\right)}{\mathbb{E}}\left[p^{\star}(y\succ y^{\prime}|x)\log\sigma\left(\hat{r}_{\theta}(x,y)-\hat{r}_{\theta}(x,y^{\prime})\right)+p^{\star}(y^{\prime}\succ y|x)\log\sigma\left(\hat{r}_{\theta}(x,y^{\prime})-\hat{r}_{\theta}(x,y)\right)\right]\;.$
@@ 3 Exact Optimization: Fine-grained Performance Gap Induced by Model Mis-specification
We analyze the behavior of RLHF and DPO in the idealized setting of exact optimization, where both methods have access to infinite preference data and can optimize their respective objectives without statistical or computational error. Our focus is on the performance gap induced by the representation gap – that is, the difference between the true optimal policy and the best policy each method can produce, as determined by the expressiveness of the reward and policy model classes.
Recall that $r_{\textup{RLHF}}\in\mathcal{F}$ is the solution computed by exact optimization of reward learning, $\pi_{\textup{RLHF}}\in\Pi$ is the solution computed by exact optimization of policy learning given $r_{\textup{RLHF}}$, and $\pi_{\textup{DPO}}\in\Pi$ is the solution computed by exact optimization of DPO.
3.1 No Model Mis-specification
We begin with the fully realizable setting, where both the ground-truth reward function and the optimal policy lie within their respective model classes. While this assumption is often unrealistic in practice, it serves as a clean baseline and has been the main focus of most prior theoretical analyses (Xiong et al., 2024; Shi et al., 2025; Feng et al., 2025; Swamy et al., 2025).
@@ 3.1 No Model Mis-specification
Both RLHF and DPO are capable of recovering the true optimal policy under ideal conditions. In this regime, RLHF directly optimizes $V_{r^{\star}}^{\pi_{\theta}}$ in the policy learning stage. Proof deferred to Section B.1.
@@
Under Condition 1, $V_{r^{\star}}^{\pi_{\textup{RLHF}}}=V_{r^{\star}}^{\pi_{\textup{DPO}}}=V_{\Pi}^{\star}$.
@@
with a ratio $\alpha_{1}=1$ and $\alpha_{2}=\underset{y,y^{\prime}\sim\pi_{\theta}}{\mathbb{E}}\exp(\hat{r}_{\theta}(x,y)-\hat{r}_{\theta}(x,y^{\prime}))\;$.
@@
$\mathcal{L}_{\textup{DPO}}^{\textup{online}}(\pi_{\theta}(\cdot|x)){\stackrel{{\scriptstyle\scriptstyle\nabla}}{{=}}}\frac{2\beta}{\mathsf{sg}\left(Z_{\theta}(x)\right)}$ $\Bigg\{-\left[\underset{y\sim\pi_{\theta}(\cdot|x)}{\mathbb{E}}[r(x,y)]-\beta\mathsf{KL}\left(\pi_{\theta}(\cdot|x)\|\pi_{{\mathsf{ref}}}(\cdot|x)\right)\right]$
$+\frac{1}{4\beta}\underset{y,y^{\prime}\sim\mathsf{sg}\left(\pi_{\theta}(\cdot|x)\right)}{\mathbb{E}}\left[\epsilon_{y,y^{\prime}}\cdot\left[\left(r^{\star}(x,y)-r^{\star}(x,y^{\prime})\right)-\left(\hat{r}_{\theta}(x,y)-\hat{r}_{\theta}(x,y^{\prime})\right)\right]^{2}\right]\Bigg\}\;,$
where $\epsilon_{y,y^{\prime}}\in\mathbb{R}$ are noises s.t. $|\epsilon_{y,y^{\prime}}|\leq\frac{\delta}{6\sqrt{3}\sigma^{\prime(}R_{\max}+\delta)}$ and $Z_{\theta}(x):=\underset{y,y^{\prime}\sim\pi_{\theta}(\cdot|x)}{\mathbb{E}}1/\sigma^{\prime}(\hat{r}_{\theta}(x,y)-\hat{r}_{\theta}(x,y^{\prime}))$ can be viewed as adaptive step sizes for different prompts.
@@
This result indicates that, with an appropriate sampler, the objective of online DPO can approximate the true value function in prompt level. However, the second-order deviation can become substantial when $R_{\max}$ is large, or the ground-truth reward is poorly fitted. In such scenarios, the objective of online DPO may significantly deviate from the value function, leading to degraded convergence or even divergence. Proof deferred to Section B.2.
3.2 Policy Model Mis-specification
We now examine the setting where the ground-truth reward function is realizable ($r^{\star}\in\mathcal{F}$), but the optimal policy is non-realizable by the policy class ($\pi^{\star}\notin\Pi$). This case can be referred to Nika et al. (2024), who point out that the optimal policy could be more complicated than the optimal reward, and Swamy et al. (2025), who attribute this scenario to generation-verification gaps in fine-tuning.
@@ 3.2 Policy Model Mis-specification
In this case, RLHF has a structural advantage: it can recover the exact reward and then compute the best possible policy within $\Pi$. In contrast, DPO bypasses reward modeling and directly learns a policy from preferences, which may lead to sub-optimal behavior due to mismatches between preference-based objectives and reward-based value functions. The following proposition provides a concrete example where DPO fails to recover the best achievable policy, even under exact optimization. Proof deferred to Section B.3.
@@
Under Condition 2, $V_{\Pi}^{\star}=V_{r^{\star}}^{\pi_{\textup{RLHF}}}\geq V_{r^{\star}}^{\pi_{\textup{DPO}}}$, and there exists an environment s.t. $V_{r^{\star}}^{\pi_{\textup{RLHF}}}>V_{r^{\star}}^{\pi_{\textup{DPO}}}$.
Furthermore, we show that online DPO cannot close this gap, even when equipped with PILAF sampler. A numerical proof is deferred to Section B.8.
@@
Our key insight is that a strict performance gap between RLHF and DPO can exist under policy model mis-specification, and importantly, even sophisticated samplers like PILAF may fail to close the gap, an important nuance that, to our knowledge, has been overlooked in prior studies.
3.3 Reward Model Mis-specification
We now consider the setting where the ground-truth reward function $r^{\star}$ is not realizable by the reward model class $\mathcal{F}$, while the optimal policy $\pi^{\star}$ lies within the policy class $\Pi$. As discussed in Swamy et al. (2024), two-stage RLHF can only lose information during reward learning, which will be highlighted under reward model mis-specification.
@@ 3.3 Reward Model Mis-specification
In this setting, RLHF is vulnerable to reward mis-specification: the learned mis-specified reward model $r_{\textup{RLHF}}$ could significantly deviate from the ground-truth reward $r^{\star}$, causing the subsequent policy optimization to yield a sub-optimal solution even though $\pi^{\star}\in\Pi$. Conversely, DPO has a clear advantage: it can directly fit a policy to the observed preference data and thus recover $\pi^{\star}$ without incurring reward modeling error. Proof deferred to Section B.4.

### v3

@@ Abstract
We present a fine-grained theoretical analysis of the performance gap between reinforcement learning from human feedback (RLHF) and direct preference optimization (DPO) under a representation gap. Our study decomposes this gap into two sources: an explicit representation gap under exact optimization and an implicit representation gap under finite samples. In the exact optimization setting, we characterize how the relative capacities of the reward and policy model classes influence the final policy qualities. We show that RLHF, DPO, or online DPO can outperform one another depending on type of model mis-specifications. Notably, online DPO can outperform both RLHF and standard DPO when the reward and policy model classes are isomorphic and both mis-specified. In the approximate optimization setting, we provide a concrete construction where the ground-truth reward is implicitly sparse and show that RLHF requires significantly fewer samples than DPO to recover an effective reward model, highlighting a statistical advantage of two-stage learning. Together, these results provide a comprehensive understanding of the performance gap between RLHF and DPO under various settings, and offer practical insights into when each method is preferred.
@@ 1 Introduction
However, the assumptions of tabular parameterization and realizability often do not hold in practice, particularly when the reward model is significantly smaller than the policy model (e.g., 66B vs. 175175B in Ouyang et al. (2022), indicating a clear disparity in representational capacity), when the policy model class is heavily restricted due to limited computational resources, or when the reward model is sub-optimal owing to limited preference data. These situations are examples of representation gap, a common issue in practice due to limitations in model capacity or data. Consequently, one should not expect DPO to perform identically to RLHF. This motivates the central question of our investigation:
@@
To quantify the problem, we choose the performance metric as the expected value of the original regularized bandit problem using the ground-truth reward $r^{\star}$ ($x$ is a prompt, and $y$ is a response):
$V_{r^{\star}}^{\pi}:=\mathbb{E}_{x\sim\rho}\left[\mathbb{E}_{y\sim\pi(\cdot|x)}[r^{\star}(x,y)]-\beta\mathsf{KL}\left(\pi(\cdot|x)\|\pi_{{\mathsf{ref}}}(\cdot|x)\right)\right]\;,$
where $\rho$ is a pre-fixed distribution over prompts, $\pi$ is a distribution over responses given prompts, and $\pi_{{\mathsf{ref}}}$ is a fixed reference policy. Let $\pi^{\star}:=\operatorname{argmax}_{\pi}V_{r^{\star}}^{\pi}$ be the ideal optimal policy.
Our contributions. We study the performance differences between RLHF and DPO, from optimization and statistic perspectives. Our contributions are listed as follows:
$V_{r^{\star}}^{\pi_{\textup{RLHF}}}=V_{r^{\star}}^{\pi_{\textup{DPO}}}=V_{r^{\star}}^{\pi^{\star}}$ $\mathcal{L}_{\textup{DPO}}^{\textup{online}}(\pi(\cdot|x))\propto-V_{r^{\star}}^{\pi(\cdot|x)}+\mathcal{O}(\delta^{2})$ $V_{r^{\star}}^{\pi_{\textup{RLHF}}}\leq V_{r^{\star}}^{\pi_{\textup{DPO}}}$ $V_{r^{\star}}^{\pi_{\textup{RLHF}}}\geq V_{r^{\star}}^{\pi_{\textup{DPO}}}$ $V_{r^{\star}}^{\pi_{\textup{RLHF}}}\geq V_{r^{\star}}^{\pi_{\textup{DPO}}^{\textup{online}}}$ It depends on
@@
$V_{r^{\star}}^{\pi_{\textup{RLHF}}}=V_{r^{\star}}^{\pi_{\textup{DPO}}}$$V_{r^{\star}}^{\pi_{\textup{RLHF}}}\leq V_{r^{\star}}^{\pi_{\textup{DPO}}^{\textup{online}}}$No Model Mis-specificationPolicy Model Mis-specificationReward Model Mis-specificationIsomorphic Model Mis-specification
@@
$\bullet$ When assuming exact optimization, i.e., optimization with infinite data, we remove the widely-used realizability assumption, and study the fine-grained performance gap under different settings of model mis-specifications in Section 3. Main results are visualized in Figure 1.
No model mis-specification: RLHF and DPO policies both achieve the performance of $\pi^{\star}$, and online DPO can further close the gap between optimization paths.
Policy model mis-specification: RLHF policy is still optimal under the model class, while the DPO policy can be sub-optimal, and online DPO cannot bridge the gap.
Reward model mis-specification: DPO policy is still optimal, while the RLHF policy can be sub-optimal due to learning based on a sub-optimal reward model.
Double model mis-specification: When policy and reward model classes are isomorphic, they should have identical performance, while online DPO can outperform both of them. Otherwise, there is no consistent performance gap, and the comparison result depends on the qualities of (surrogate) reward models.
$\bullet$ For approximate optimization, i.e., the finite-sample regime, we study the performance gap incurred by statistical efficiencies in Section 4. We show that DPO training can distort the intrinsic structure of the reward, and demonstrate this effect under a linear parameterization. We then construct a simple task where the ground-truth reward feature has dimension $d$ and sparsity $k$, and the total number of samples is $n$. Even without mis-specifications, we can reveal a separation between RLHF and DPO under this setting: the estimation error of DPO is $\Omega(d/n)$, while reward learning in RLHF can effectively leverage sparsity, reducing the error to $\tilde{\mathcal{O}}(\sqrt{k\log d/n})$. This result indicates that DPO is less data-efficient than RLHF, leading to inferior performance.
Finally, we conduct controlled experiments to corroborate these theoretical findings and connect existing empirical studies with our theoretical claims in Section 5.
@@ 2 Preliminaries
Notation. Let $\sigma:\mathbb{R}\rightarrow\mathbb{R}$ be the sigmoid function, where $\sigma(x)=1/(1+\exp(-x))$. For any set $\mathcal{X}$, $\Delta(\mathcal{X})$ represents the set of probability distributions over $\mathcal{X}$. $\mathsf{sg}\left(\right)$ is the stopping-gradient operator, where $\nabla_{\theta}[\mathsf{sg}\left(f(\theta)\right)]=\boldsymbol{0}$. Let $e_{k}$ be a one-hot vector with $1$ on its $k\textsuperscript{th}$ entry and 0 on other entries. For any vector $x$, let $x_{k}$ be its $k\textsuperscript{th}$ entry. We use $f(\theta){\stackrel{{\scriptstyle\scriptstyle\nabla}}{{=}}}g(\theta)$ to indicate $\nabla_{\theta}f(\theta)=\nabla_{\theta}g(\theta)$. We define the difference operator as $\Delta_{x_{1},x_{2}}(f)=f(x_{1})-f_{1}(x_{2})$.
Policy space. There is a prompt space $\mathcal{X}$, a response space $\mathcal{Y}$, and a reward function $r:\mathcal{X}\times\mathcal{Y}\to\mathbb{R}$. A policy $\pi:\mathcal{X}\rightarrow\Delta(\mathcal{Y})$ represents a probability distribution over responses given a prompt. Note that, we sometimes omit the prompt $x$ for simplicity, so that $\pi\in\Delta(\mathcal{Y})$.
@@
$V_{r}^{\pi(\cdot|x)}$ $:=\left[\underset{y\sim\pi(\cdot|x)}{\mathbb{E}}[r(x,y)]-\beta\,\mathsf{KL}\left(\pi(\cdot|x)\|\pi_{{\mathsf{ref}}}(\cdot|x)\right)\right]\;,$
$V_{r}^{\pi}$ $:=\underset{x\sim\rho}{\mathbb{E}}V_{r}^{\pi(\cdot|x)}\;,$
where $\beta>0$ is the regularization coefficient, $\rho\in\Delta(\mathcal{X})$ is a pre-fixed distribution over prompts, and $\pi_{{\mathsf{ref}}}$ is a fixed reference policy. Let $r^{\star}$ denote the ground-truth reward function, and $\pi^{\star}$ denote the optimal policy for $V_{r^{\star}}^{\pi}$. A well-known fact (Rafailov et al., 2023) is that $\pi^{\star}(y|x)=\pi_{{\mathsf{ref}}}(y|x)\exp(r^{\star}(x,y)/\beta)/Z(x)$, where $Z(x):=\sum_{y\in\mathcal{Y}}\pi_{{\mathsf{ref}}}(y|x)\exp(r^{\star}(x,y)/\beta)$ is the partition function. The goal of preference-based policy learning is to find a policy $\pi_{\theta}\in\Pi$ that maximizes $V_{r^{\star}}^{\pi_{\theta}}$. We define the oracle value as $V^{\Pi}_{r^{\star}}:=\max_{\pi\in\Pi}V_{r^{\star}}^{\pi}$.
@@
Empirical preference dataset. In practice, people first collect a pair dataset $\mathcal{D}^{\dagger}=\{x^{(i)},y_{1}^{(i)},y_{2}^{(i)}\}_{i=1}^{n}$, and then ask human annotators to label these pairs to get a human preference dataset $\mathcal{D}=\{x^{(i)},y_{w}^{(i)},y_{l}^{(i)}\}_{i=1}^{n}$. Following BT model, $y_{1}^{(i)}$ is preferred over $y_{2}^{(i)}$ given prompt $x^{(i)}$, (i.e. $y_{w}=y_{1}$ and $y_{l}=y_{2}$), w.p. $p^{\star}(y_{1}^{(i)}\succ y_{2}^{(i)}|x^{(i)})$.
Two-stage approach of RLHF. RLHF proceeds in two stages. First, the reward learning stage finds a reward model $r_{\textup{RLHF}}\in\mathcal{F}$ by maximizing the MLE objective:
$r_{\textup{RLHF}}=\underset{r_{\phi}\in\mathcal{F}}{\operatorname{argmax}}\;\underset{x\sim\rho;y,y^{\prime}\sim\pi_{{\mathsf{ref}}}(\cdot|x)}{\mathbb{E}}\sum_{\{y_{1},y_{2}\}=\{y,y^{\prime}\}}p^{\star}(y_{1}\succ y_{2}|x)\log\sigma(r_{\phi}(x,y_{1})-r_{\phi}(x,y_{2}))\;.$
And for approximate optimization, $r_{\text{RLHF}}$ is estimated from a finite human preference dataset. Then using the reward model $r_{\textup{RLHF}}$, the policy learning stage returns $\pi_{\textup{RLHF}}=\operatorname{argmax}_{\pi\in\Pi}\;V_{r_{\textrm{RLHF}}}^{\pi}$.
Direct approach of DPO. By leveraging the surrogate reward $\hat{r}_{\theta}(x,y):=\beta\log\frac{\pi_{\theta}(x,y)}{\pi_{{\mathsf{ref}}}(x,y)}$, DPO bypasses reward learning and directly learns policy from preference:
$\pi_{\text{DPO}}=\underset{\pi_{\theta}\in\Pi}{\operatorname{argmax}}\;\underset{x\sim\rho;y,y^{\prime}\sim\pi_{{\mathsf{ref}}}(\cdot|x)}{\mathbb{E}}\sum_{\{y_{1},y_{2}\}=\{y,y^{\prime}\}}p^{\star}(y_{1}\succ y_{2}|x)\log\sigma\left(\hat{r}_{\theta}(x,y_{1})-\hat{r}_{\theta}(x,y_{2})\right)\;.$
For approximate optimization, $\pi_{\text{DPO}}$ is estimated from an empirical preference dataset. We also consider an online variant of DPO (Xiong et al., 2024), where the pairwise data are sampled from a distribution $\pi^{\mathsf{s}}$ which could depend on the current policy. It then minimizes the modified loss:
$\mathcal{L}_{\textup{DPO}}^{\textup{online}}(\pi_{\theta}(\cdot|x))=-\underset{y,y^{\prime}\sim\mathsf{sg}\left(\pi^{\mathsf{s}}(\cdot|x)\right)}{\mathbb{E}}\sum_{\{y_{1},y_{2}\}=\{y,y^{\prime}\}}p^{\star}(y_{1}\succ y_{2}|x)\log\sigma\left(\hat{r}_{\theta}(x,y_{1})-\hat{r}_{\theta}(x,y_{2})\right)\;.$
@@ 3 Exact Optimization: Fine-grained Performance Gap Induced by Model Mis-specification
We analyze the behavior of RLHF and DPO in the idealized setting of exact optimization, where both methods have access to infinite preference data and can optimize their respective objectives without statistical or computational error. Recall that $r_{\textup{RLHF}}\in\mathcal{F}$ is the solution computed by exact optimization of reward learning, $\pi_{\textup{RLHF}}\in\Pi$ is the solution computed by exact optimization of policy learning given $r_{\textup{RLHF}}$, and $\pi_{\textup{DPO}}\in\Pi$ is the solution computed by exact optimization of DPO. We can bound the sub-optimality of each algorithm using the mis-specification error (see calculations in Appendix B.12), but in this section our focus is on the performance gap induced by model mis-specification, that is, the difference between the best policy each method can produce, as determined by the expressiveness of the reward and policy model classes.
No Model Mis-specification. We begin with the fully realizable setting, where both the ground-truth reward function and the optimal policy lie within their respective model classes. While this assumption is often unrealistic in practice, it serves as a clean baseline and has been the main focus of most prior theoretical analyses (Xiong et al., 2024; Shi et al., 2025; Feng et al., 2025; Swamy et al., 2025).
@@
Both RLHF and DPO are capable of recovering the true optimal policy under ideal conditions. In this regime, RLHF directly optimizes $V_{r^{\star}}^{\pi_{\theta}}$ in the policy learning stage. Proof deferred to Appendix B.1.
@@
Under Condition 1, $V_{r^{\star}}^{\pi_{\textup{RLHF}}}=V_{r^{\star}}^{\pi_{\textup{DPO}}}=V^{\Pi}_{r^{\star}}$.
@@
with ratio $\alpha_{1}=1,\alpha_{2}=\underset{y,y^{\prime}\sim\pi_{\theta}}{\mathbb{E}}\exp(\hat{r}_{\theta}(x,y)-\hat{r}_{\theta}(x,y^{\prime}))\;$.
@@
$\mathcal{L}_{\textup{DPO}}^{\textup{online}}(\pi_{\theta}(\cdot|x)){\stackrel{{\scriptstyle\scriptstyle\nabla}}{{=}}}\frac{2\beta}{\mathsf{sg}\left(Z_{\theta}(x)\right)}\Bigg\{-V_{r}^{\pi(\cdot|x)}+\frac{1}{4\beta}\underset{y,y^{\prime}\sim\mathsf{sg}\left(\pi_{\theta}(\cdot|x)\right)}{\mathbb{E}}\left[\epsilon_{y,y^{\prime}}\cdot\Delta^{2}_{(x,y),(x,y^{\prime})}(r^{\star}-\hat{r}_{\theta})\right]\Bigg\}\;,$
where $\epsilon_{y,y^{\prime}}\in\mathbb{R}$ are noises s.t. $|\epsilon_{y,y^{\prime}}|\leq\frac{\delta}{6\sqrt{3}\sigma^{\prime(}R_{\max}+\delta)}$, $\Delta(\cdot)$ is the difference operator as defined in Section 2, and $Z_{\theta}(x):=\underset{y,y^{\prime}\sim\pi_{\theta}(\cdot|x)}{\mathbb{E}}1/\sigma^{\prime}(\hat{r}_{\theta}(x,y)-\hat{r}_{\theta}(x,y^{\prime}))$ can be viewed as adaptive step sizes for different prompts.
@@
This result indicates that, with an appropriate sampler, the objective of online DPO can approximate the true value function in prompt level. However, the second-order deviation can become substantial when $R_{\max}$ is large, or the ground-truth reward is poorly fitted. In such scenarios, the objective of online DPO may significantly deviate from the value function, leading to degraded convergence or even divergence. Proof deferred to Appendix B.2.
Policy Model Mis-specification. We now examine the setting where the ground-truth reward function is realizable ($r^{\star}\in\mathcal{F}$), but the optimal policy is non-realizable by the policy class ($\pi^{\star}\notin\Pi$). This case can be referred to Nika et al. (2024), who point out that the optimal policy could be more complicated than the optimal reward, and Swamy et al. (2025), who attribute this scenario to generation-verification gaps in fine-tuning.
@@
In this case, RLHF has a structural advantage: it can recover the exact reward and then compute the best possible policy within $\Pi$. In contrast, DPO bypasses reward modeling and directly learns a policy from preferences, which may lead to sub-optimal behavior due to mismatches between preference-based objectives and reward-based value functions. The following proposition provides a concrete example where DPO fails to recover the best achievable policy, even under exact optimization. Proof deferred to Appendix B.3.
@@
Under Condition 2, $V^{\Pi}_{r^{\star}}=V_{r^{\star}}^{\pi_{\textup{RLHF}}}\geq V_{r^{\star}}^{\pi_{\textup{DPO}}}$, and there exists an environment s.t. $V_{r^{\star}}^{\pi_{\textup{RLHF}}}>V_{r^{\star}}^{\pi_{\textup{DPO}}}$.
Furthermore, we show that online DPO cannot close this gap, even when equipped with PILAF sampler. A numerical proof is deferred to Appendix B.8.
@@
Our key insight is that a strict performance gap between RLHF and DPO can exist under policy model mis-specification, and importantly, even sophisticated samplers like PILAF may fail to close the gap. This is an important nuance that has been overlooked in prior studies.
Reward Model Mis-specification. We now consider the setting where the ground-truth reward function $r^{\star}$ is not realizable by the reward model class $\mathcal{F}$, while the optimal policy $\pi^{\star}$ lies within the policy class $\Pi$. As discussed in Swamy et al. (2024), two-stage RLHF can only lose information during reward learning, which will be highlighted under reward model mis-specification.
@@
In this setting, RLHF is vulnerable to reward mis-specification: the learned mis-specified reward model $r_{\textup{RLHF}}$ could significantly deviate from the ground-truth reward $r^{\star}$, causing the subsequent policy optimization to yield a sub-optimal solution even though $\pi^{\star}\in\Pi$. Conversely, DPO has a clear advantage: it can directly fit a policy to the observed preference data and thus recover $\pi^{\star}$ without incurring reward modeling error. Proof deferred to Appendix B.4.